OpenStudy (anonymous):

Which literal equations are equivalent to K= F/x? Choose all answers that are correct. A. F=kx B. X=kF C. F=k/x D. x=F/k @iGreen @dan815 @Texas_Girl @Troy. @Blu-Girl @baking4lyf @animal_lover36 @Firefly007 @geerky42 @just_one_last_goodbye @KittyT @linn99123 @mathmath333 Fan and Medal Please help!!!

OpenStudy (midhun.madhu1987):



OpenStudy (igreen):

@midhun.madhu1987 Don't you mean multiply both sides by \(x\)?

OpenStudy (midhun.madhu1987):

yep... sorry... my mistake :(

OpenStudy (igreen):

\(k = \dfrac{F}{x}\) \(kx = \dfrac{F}{x} \cdot x\) \(kx = \dfrac{F}{\bcancel x} \cdot \bcancel x\)

OpenStudy (anonymous):

im so so sorry but i suck at math lol

OpenStudy (anonymous):

I'm completely and utterly lost

OpenStudy (anonymous):

math is (Mental, Abuse, To, Humans)

OpenStudy (igreen):

The \(x\)'s cancel out, leaving you with what? @WolvesAllTheWay2002 \(kx = \dfrac{F}{\bcancel x} \cdot \bcancel x\)

OpenStudy (anonymous):

k/x=F?

OpenStudy (igreen):

No, you keep the \(kx\).

OpenStudy (anonymous):

okay so that would be what

OpenStudy (igreen):

\(kx = F\)

OpenStudy (anonymous):

oh its A.?

OpenStudy (igreen):

That's 1 of the chocies.

OpenStudy (igreen):

*answers

OpenStudy (anonymous):

and B?

OpenStudy (igreen):

There's another one.. \(kx = F\) Now let's solve for \(x\). We divide \(x\) to both sides: \(\dfrac{kx}{x} = \dfrac{F}{x}\) The \(x\)'s cancel out, leaving you with what? \(\dfrac{kx}{\bcancel x} = \dfrac{F}{\bcancel x}\)

OpenStudy (anonymous):

how about c?

OpenStudy (igreen):

Sorry, we divide k to both sides, which gives us: \(x = \dfrac{F}{k}\)

OpenStudy (igreen):

So what are your two answers going to be? @WolvesAllTheWay2002

OpenStudy (anonymous):

A,B